If there is a mass percentage of 24.5% gold in an ore vein. How many moles of gold can be produced from 1.00 x 103 kg of ore? Assume 100% recovery


1
Expert's answer
2021-10-30T01:11:06-0400

Proportion:

1000 kg – 100 %

x kg – 24.5 %

"x = \frac{1000 \times 24.5}{100} = 245 \; kg = 245 \times 10^3 \; g"

MM(Au) = 196.967 g/mol

"n(Au) = \frac{245 \times 10^3}{196.967} = 1.243 \times 10^3 \;mol = 1243 \;mol"

Answer: 1243 mol